# Mobile Edge Generation-Enabled Digital Twin: Architecture Design and Research Opportunities

### v1

@@ III-B Convergence between MEG and DT
  • MEG Aided DT: MEG enables autonomous creation, maintenance, and validation of DT on mobile devices. Specifically, the decentralized GAI sub-models can be trained to minimize the divergence between the digital model distribution and the real target distribution, thus accurately modelling the real-world complex systems to construct high-fidelity digital replications. Notably, MEG can creatively generate vivid images, videos, and voices in the virtual world by synthesizing the inputs from different mobile UEs. Furthermore, MEG can learn to emulate the expert behaviours in the high-fidelity virtual world for decision-making. Additionally, testing scenarios can be generated via MEG for validation, thus assisting the online self-prediction, self-diagnose, and self-management of DT systems.
  • DT Aided MEG: A virtual twin can be created by DT to capture the dynamics of mobile edge networks based on data analytics and multi-physics simulation. This can enable self-optimized configuration of MEG. Specifically, DT can simulate the real-world channel models and track UEs’ mobility, thus empowering low-overhead and high-reliability sketch/seed transmissions for MEG. Furthermore, based on the predictive analytics of UEs’ behaviours and environmental changes, reliable and secure distributed computing schemes for MEG can be derived in DT, which balances the computation/communication workloads and diagnoses potential malicious attacks of mobile devices.
@@ IV Case Studies
In this section, we identify suitable generation mechanisms/protocols for multi-modal DT generation and demonstrate the performance of E2U seed-based image generation via numerical examples.
IV-A Multi-Modal Generation
DT involves generations of multi-modal contents, including text, images, videos, and interactive scenes, etc. We summarize suitable protocols for different generation tasks in Fig. 5. Text generation can be performed via centralized GAI models as it requires only a small transmission bandwidth. Here, we discuss image/video generation and interactive scene generation.
  • Image Generation: Conventional image generation tasks often perform generation based on prompts. The prompts can be text, image, or even multi-modal data. Regarding text-to-image generation, the bandwidth required for text transmission is much less than that of image. Hence, to reduce the communication overheads, it is suitable to employ the E2U/U2E mechanism for text-to-image generation at UEs/ESs. Regarding image-to-image generation, SEU/PEU mechanism can be employed for sequential or synergetic feature encoding. On the other hand, both seeds and sketches are feasible for conveying compressed information, but have distinct advantages. The seed can achieve higher compression rates, while sketch can provide human-readable visual previews.
  • Video Generation: The video transmission requires enormous bandwidth, especially when high-resolution virtual worlds should be constructed. For image-to-video generation, applying SEU protocol for video generation can efficiently reduce the size of the video file. Nevertheless, the users can only receive the generated video after the generation process, which may increase the waiting time for users. In this case, the PEU generation has a chance to provide a simultaneous generation and presentation of the video. To enhance compression ratio, seed-based generation is preferable in this scenario.
  • Interactive Scene: A key function of DT is to provide interaction. Thus, the generation of interactive scenes is a crucial part of MEG-DT. Compared with image/video generation, interactive scenes have higher requirements for latency. Therefore, the PEU protocol should be adopted in the generation of interactive scenes for virtual-reality convergence. In addition, as interactive scenes require continuous visualization results (even with insufficient clarity), sketch is more suitable for serving the generation of interactive scenes.
(a) Generation quality.
@@ IV-A Multi-Modal Generation
Figure 6: Performance comparisons for generation approaches.
@@ IV-B Performance Evaluation
In this part, we provide case studies of text-to-image generation for the proposed MEG-DT architecture. The seed-based E2U generation protocol is investigated under the compact additive white Gaussian noise channel. To train distributed GAI sub-models for seed encoding/decoding, images generated by centralized GAI and transmitted under perfect channel conditions are set as the ground truth. Given the ground truth, we calculate the Fréchet Inception Distance (FID) score to measure the distortion of reconstructed images via MEG-DT. We assume $3232$-bit floating point number for transmissions. The following three generation approaches are considered in the simulation:
  • Centralized Generation: A pretrained Stable Diffusion model [7] is utilized as baseline for centralized generation. An image with $33.5533.55$ Mbits is generated by ES and is transmitted to UE in the downlink.
  • MEG-DT: By splitting the pretrained Stable Diffusion model, a denoising diffusion implicit models (DDIM) sampler is deployed on ES, which samples a $0.520.52$-bit latent feature as a seed for distributed generation on UE.
  • E2E-MEG-DT: ES and UE train GAI sub-models in noisy wireless channels to achieve end-to-end coding and transmission of $0.260.26$-bit seeds. A JSCC encoder/decoder is integrated into Stable Diffusion model for mapping seeds’ source signals into channel symbols.
Fig. 6 presents the numerical results, where a $1024\times 10241024 × 1024$ flower scene is generated by text prompt in the virtual twin. By transmitting seeds instead of images through the noisy wireless channels, MEG-DT schemes can achieve noise-resistant generation compared to centralized generation, whilst significantly reducing overheads. Furthermore, the E2E-MEG-DT scheme is more robust and exhibits a significantly higher gains in both generation quality and FID score, even in low SNR conditions. It is also worth noting that under high SNR conditions (SNR>20absent20>20> 20), additional slight distortion is introduced due to the difficulty in perfectly training the end-to-end encoder and decoder.
@@ V-A1 Smart City
: Smart city involves densely deployed sensors, interconnected vehicles/robots, social networks, and cellular/WiFi networks. With the huge amounts of data continuously produced by these entities, virtual city twins can be constructed via MEG-DT to integrate observations of massive mobile devices. This will empower accurate time-series traffic forecasting, social event forecasting, and rational decision-making for city governors, traffic administrator, and network operators, etc.
@@
: Industry 4.0 requires cyber-physical integration for smart manufacturing [13]. MEG can easily build cyber models of products using prompts and sensing data, thus achieving personalized and low-cost product designs. During manufacturing process, on-device virtual twins of smart factories can be created, thus supporting 24/7 real-time monitoring and quality control.
V-A3 Extended Reality (XR) and Metaverse
: The distributed GAI sub-models can synthesize diverse datasets collected by different devices, thus realizing the convergence between virtual and real worlds to build XR and metaverse applications. For instance, the vehicular XR metaverse can be built by synthesizing massive traffic and driving data, thus improving robustness, driving safety, traffic control efficiency, and sustainability [14].
@@ V-A4 Healthcare
: Enabling mobile AIGC in DT is promising to provide personalized healthcare [15]. MEG-DT delivers a low-latency distributed solution for mobile AIGC, which can simulate data of rare diseases, establish testbeds for comprehensive drug validations and analytics, and provide ubiquitous personalized medical services.
@@ V-B Open Problems of MEG-DT
  • Security: Although seed-based generation enables privacy preservation, security remains an open challenge in MEG-DT especially for sensitive applications. Specifically, when the sketches/seeds are attacked or corrupted by poisoning, the content generation process may be collapsed. Furthermore, since sketches and seeds generally contain important object features and substantial semantic information, some privacy-sensitive information (e.g., local sensing data) may be eavesdropped by attackers through reverse engineering techniques, leading to privacy disclosure. These vulnerabilities pose an urgent requirement to incorporate advanced security mechanisms into the MEG-DT architecture, such as differential privacy and blockchain techniques.
@@
  • Scalable Learning and Compression: There are several candidate techniques for the compression of sketch/seed and GAI sub-models [12]. Specifically, to achieve a compact model representation, inessential components can be removed by pruning, network sparsification, and knowledge distillation algorithms, thus shrinking the large-scale models with achievable performance loss. Another approach is to decompose cumbersome learning parameters into small-sized tensors via tensor decomposition, which can relieve memory requirements and computational complexity. Instead of reducing the number of model parameters, the quantization method focuses on achieving low-bit features. Note that the compression and transmission of learning features on UEs are simultaneously subject to the hardware complexity limitation and the radio resource scarcity. This constitutes a research interest of developing scalable learning and compression techniques that can balance between compression ratio, hardware applicability, and transmission quality (e.g., sketch/seed distortion, latency and reliability).
  • Distributed Computing and Coordination: MEG-DT depends on efficient distributed computing and coordination. On the one hand, the distributed deployment of GAI sub-models requires adaptive configuration to achieve the optimal performance with acceptable latency and complexity. For individual generation, the combined use of sketches and seeds should be designed to achieve extreme generation quality. For coordinated generation, UEs should dynamically determine the optimal set of interacting devices and the optimal frequency for sketch/seed interactions. On the other hand, the sensing, communication and computing resources over multiple devices should be coordinated to accurately track environment changes to refresh digital replica, as well as automatically making decisions in response to both current conditions and predictive changes. This requires the coordination between heterogeneous hardware devices and functionalities.
@@
  • Personalization and Customization: Personalization is a crucial aspect for MEG-DT to meet users’ individual expectations. For personalized generation, GAI sub-models need to be customized on individual UEs that experience different data distributions, performance metric requirements, and hardware/software limitations. The personalized customization can be mainly achieved via two approaches, i.e., data augmentation and model adaptation. On the one hand, data augmentation directly tackles the data distribution drift, which pretrains the globally shared GAI sub-models with heterogeneous data and then fine-tunes the pretrained sub-models via UE’s local data. On the other hand, model adaptation achieves enhanced and accelerated personalization of GAI sub-models via few-shot learning (e.g., meta-learning and transfer learning) or knowledge transfer techniques. Specifically, meta-learning learns an optimal GAI meta-model that can fast generalize across a range of tasks by maximizing the model parameter sensitivity. Based on the meta-model, personalized sub-models for particular device can be achieved via few-shot adaptation. Moreover, transfer learning focuses on diminishing the domain discrepancy between the globally shared sub-models and personalized sub-models. Instead of transferring model parameters, knowledge transfer further shares knowledge between initial sub-models and personalized sub-models via soft predictions, which enables heterogeneous sub-model structures on different UEs.

### v3

@@ III-B Convergence between MEG and DT
  • MEG Aided DT: MEG enables autonomous creation, maintenance, and validation of DT on mobile devices. Specifically, the decentralized GAI sub-models can be trained to minimize virtual-reality divergence, thus constructing high-fidelity digital replications. Moreover, Vivid images, videos, and voices can be created by MEG by synthesizing inputs from mobile UEs based on multi-UE coordination. Further, MEG can emulate expert behaviours in virtual worlds for decision-making. Then, testing scenarios can be generated for validation, thus assisting the online self-prediction, self-diagnose, and self-management of DT systems.
  • DT Aided MEG: DT can capture system dynamics of mobile edge networks based on data analytics and multi-physics simulation, thus achieving self-optimized configuration of MEG. For single-user generation, DT can empower low-overhead and high-reliability feature transmissions by simulating real-world channel environment and track UEs’ mobility. For multi-user/ES cooperative coordination, DT can provide predictive analytics of UEs’ behaviours and environmental changes and derive reliable and secure distributed computing schemes, which balances the computation/communication workloads and diagnoses potential malicious attacks for MEG.
@@ IV Case Studies
This section discusses the applicability of distributed deployment and generation strategies of GAI models in various scenarios. Further, we provide numerical examples to demonstrate the performance of E2U seed-based image generation.
IV-A Applicability Analyses
DT involves the editing, transformation, and synthesis of multi-modal contents across texts, voices, images, videos, etc. The applicabilities of distributed deployment and generation strategies are summarized in Fig. 5, as discussed as follows.
  • For small-scale/modest-scale generation tasks (e.g., text creation and text-to-image generation), the distributed deployment across a single-ES single-UE system can satisfy computational requirements for popular GAI models, such as Stable Diffusion and DALL-E. The applicable generation mechanisms and feature representations can be analyzed as follows. (i) Generation mechanisms: Different generation mechanisms can be exploited for data compressions in DL/UL transmissions on demand. For UE-side text-to-image generations, since text uploading consumes much less resources than image downloading, E2U generation can be adopted to encode image features in DL transmissions. Similarly, U2E generation can be applied in the ES-side text-to-image generation. For image-to-image and image-to-video generations, both UL and DL transmissions are bandwidth-thirsty. To reduce transmission overheads, we can exploit SEU generation to sequentially encode vision features at the UE and the ES. Moreover, PEU generation enables low-latency synthesis of images/videos/musics from the ES and the UE. (ii) Feature representations: For feature representations, sketch-based generation enables faster content selections and inspections. Before reconstructing high-dimension contents, users can quickly glance at sketches to filter out undesirable contents. By contrast, the seed-based generation can achieve higher compression rates especially for video streams and preserve data privacy.
  • For large-scale generation tasks (such as ultra-high-resolution (UHD) images/videos generation and metaverse applications), the required GPU memory and computing power may overwhelm the capacities of a single-ES single-UE system. A tractable solution is to partition large-scale contents into multiple data slices that can be cooperatively generated by multiple interactive ESs/UEs. To update the local data slices at UEs, cooperative multi-user generation mechanisms (as discussed in Section II-C) can be utilized to enable low-cost synchronization and interactions. Moreover, to generate large-scale contents in parallel, different slices can be assigned to nearby ESs for cooperative multi-ES generation, thus fully utilizing edge computing resources.
(a) Scene generation quality.
@@ IV-A Applicability Analyses
Figure 6: Performance comparisons of different schemes, where various virtual scenes are created for user’s avatar in MEG-DT.
@@ IV-B Performance Evaluation
In this part, we provide case studies of text-to-image generation for the proposed MEG-DT architecture. We investigate the seed-based E2U generation under compact additive white Gaussian noise channels. We evaluate the distortion between the reconstructed images and the reference images by Fréchet inception distance (FID) scores. Here, the reference images are generated by centralized GAI and are transmitted under perfect channel conditions. We assume $1616$-bit floating point number and the average data transmission speed is $11$ Mbps. The following three generation approaches are considered:
  • Centralized Generation: The ES generates $1024\times 10241024 × 1024$ images leveraging a pre-trained Stable Diffusion XL (SDXL) model [7], which are transmitted to the UE.
  • MEG-DT: The SDXL model is split and fine-tuned over noisy channels. The latent feature is sampled by the VAE encoder and U-Net denoiser deployed at the ES, which is transmitted to the UE for image decoding.
  • E2E-MEG-DT: The latent features of MEG-DT are further compressed by merging similar neurons. A pair of JSCC encoder/decoder is jointly trained with VAE decoder, which minimizes the end-to-end image distortions over noisy wireless channels.
[…]
@@ V-A1 Smart City
Smart city involves densely deployed sensors, interconnected vehicles/robots, social networks, and cellular/WiFi networks. With the huge amounts of data continuously produced by these entities, virtual city twins can be constructed via MEG-DT to integrate observations of massive mobile devices. This will empower accurate time-series traffic forecasting, social event forecasting, and rational decision-making for city governors, traffic administrator, and network operators, etc.
@@
: Industry 4.0 requires cyber-physical integration for smart manufacturing [13]. MEG can easily build cyber-models of products using prompts and sensing data, thus achieving personalized and low-cost product designs. During manufacturing process, on-device virtual twins of smart factories can be created, thus supporting 24/7 real-time monitoring and quality control.
V-A3 Metaverse and XR
MEG-DT can synthesize diverse datasets collected by different devices, thus realizing the convergence between virtual and real worlds to build XR and metaverse applications. For instance, the vehicular XR metaverse can be built by synthesizing massive traffic and driving data, thus improving robustness, driving safety, traffic control efficiency, and sustainability [14].
@@ V-A4 Healthcare
Enabling mobile AIGC in DT is promising to provide personalized healthcare [15]. MEG-DT delivers a low-latency distributed solution for mobile AIGC, which can simulate data of rare diseases, establish testbed for comprehensive drug validations and analytics, and provide ubiquitous personalized medical services.
@@ V-B Open Problems of MEG-DT
  • Security: Although seed-based generation enables privacy preservation, security remains an open challenge in MEG-DT especially for sensitive applications. Specifically, when sketches/seeds are attacked, corrupted, or modified by viruses, the content generation process may be collapsed. This poses an urgent need for developing advanced security techniques, e.g., blockchains, for trustworthy and invulnerable distributed generation. Moreover, since sketches/seeds generally contain important object features and substantial semantic information, some privacy-sensitive information (e.g., local sensing data) may be eavesdropped by attackers through reverse engineering techniques. Hence, privacy-preserving techniques (e.g., differential privacy) can be investigated, which protects privacy by adding controllable noises to data’s statistical features.
@@
  • Scalable Learning and Compression: There are several candidate techniques for the compression of sketch/seed and GAI sub-models [12]. Specifically, to achieve compact representations, inessential components can be removed by network sparsification methods (e.g., pruning), thus shrinking redundant neurons. As visual contents may be sensitive to pruning, another approach is to decompose the high-complexity tensors into low-rank and small-sized tensors via tensor decomposition, thus relieving memory and computational overheads. Instead of reducing model parameters, the quantization method focuses on achieving low-bit features, which are suitable for low-precision computing on mobile UEs. To further mitigate distortions for various application scenarios, knowledge distillation can flexibly distil tiny GAI sub-models by re-training over dedicated datasets. In practice, MEG-DT suffers from both hardware and radio resource limitations, constituting an interest of developing scalable learning and compression techniques to balance between compression rate, hardware applicability, and generation quality.
  • Distributed Computing and Coordination: MEG-DT relies on efficient distributed computing and coordination. For individual generation, the distributed GAI sub-models need to be adaptively deployed and configured, thus enhancing the content fidelity while ensuring low latency and complexity. For coordinated generation, UEs should dynamically determine the optimal set of interacting devices and the optimal sketch/seed interaction frequency. Furthermore, to handle large-scale generation tasks, suitable data slicing and task offloading schemes are required for multi-ES cooperation. On the other hand, to accurately track environment changes whilst refreshing digital replica in the real time, the sensing, communication and computing resources over edge devices should be efficiently exploited. This requires the development of efficient coordination schemes between heterogeneous hardware devices and functionalities.
@@
  • Personalization and Customization: Personalization is a crucial aspect for MEG-DT to meet users’ individual expectations. For personalized generation, GAI sub-models need to be customized on individual UEs that experience different data distributions, performance metric requirements, and hardware/software limitations. The personalized customization can be mainly achieved via two approaches, i.e., data augmentation and model adaptation. On the one hand, data augmentation directly tackles the data distribution drift, which pre-trains the globally shared GAI sub-models with heterogeneous data and then fine-tunes the pre-trained sub-models via UE’s local data. On the other hand, model adaptation achieves enhanced and accelerated personalization of GAI sub-models via few-shot learning (e.g., meta-learning and transfer learning) or knowledge transfer techniques. Specifically, meta-learning learns an optimal GAI meta-model that can fast generalize across a range of tasks by maximizing the model parameter sensitivity. Based on the meta-model, personalized sub-models for particular device can be achieved via few-shot adaptation. Moreover, transfer learning focuses on diminishing the domain discrepancy between the globally shared sub-models and personalized sub-models. Instead of transferring model parameters, knowledge transfer further shares knowledge between initial sub-models and personalized sub-models via soft predictions, which enables heterogeneous sub-model structures on different UEs.